OpenStudy (legit_sarahh):

solve for the y value. 3x+2y=19 6x+7y=44

OpenStudy (aaliyahfairgood):

(9x+9y=63)1/9 x+y=7 y=-x+7

OpenStudy (legit_sarahh):

ok

OpenStudy (dtan5457):

@aaliyahfairgood is that your final answer?

OpenStudy (legit_sarahh):

the final answer iis y=2 i just dont know how to get it

OpenStudy (dtan5457):

Your right y does equal 2. Let me explain

OpenStudy (aaliyahfairgood):

solve for the x too? kk

OpenStudy (legit_sarahh):

ok

OpenStudy (aaliyahfairgood):

then plug x and y into both of the equations to check your answer

OpenStudy (dtan5457):

y=2 because you have the equations of 3x+2y=19 6x+7y=44 you should easily see here that the easiest form of answering this is by elimination. make the first equation have a -6 in there. you do that by multiplying that equation by -2. That makes it -6x-4y=-38 So now, your two equations are -6x-4y=-38 6x+7y=44 Add them. 3y=6 y=2

OpenStudy (aaliyahfairgood):

yep

OpenStudy (dtan5457):

TO get x, you know that y=2. Therefore, 3x+2(2)=19 That is 3x+4=19 simplify to 3x=15 x=5

OpenStudy (dtan5457):

@aaliyahfairgood 's way is actually easier if you can see what she did.

OpenStudy (aaliyahfairgood):

sorry but I am not good at explaining myself of what I did in math particularly because I think faster than I can actually verify my own answer and understand myself